A 500.0 mL sample of nitrogen is warmed from 77.0 ˚C to 86.0 ˚C. Find its new volume if the pressure remains constant.




1
Expert's answer
2022-03-20T06:38:06-0400

V1 = 500.0 mL

T1 = 77.0˚C + 273 = 350 K

P1 = P2 = constant

V2 = unknown

T2 = 86.0˚C + 273 = 359 K


Solution:

Since the pressure and amount of gas remain unchanged, Charles's law can be used.

Charles's gas law can be expressed as: V1 / T1 = V2 / T2

V1T2 = V2T1

To find the new volume, solve the equation for V2:

V2 = V1T2 / T1

V2 = (500.0 mL × 359 K) / (350 K) = 512.857 mL = 512.9 mL

V2 = 512.9 mL


Answer: The new volume is 512.9 mL